An oil drum with a diameter of 400mm and height of 2 meters is filled with water. calculate the force of pressure on the bottom of the tank.

1
Expert's answer
2022-04-11T12:06:55-0400

"p=\frac FS=\frac{mg}S=\frac{\rho Shg}S=\rho gh=20~kPa."